# On the Feasibility of Exact Unitary Transformations for Many-body Hamiltonians

### v1

@@ Abstract
Exact unitary transformations play a central role in the analysis and simulation of many-body quantum systems, yet the conditions under which they can be carried out exactly and efficiently remain incompletely understood. We show that exact transformations arise whenever the adjoint action of a unitary’s generator defines a linear map within a finite-dimensional operator space. In this regime, the Cayley–Hamilton theorem ensures the existence of a finite-degree polynomial that annihilates the adjoint map, rendering the Baker–Campbell–Hausdorff expansion finite. This perspective brings together previously disparate examples of exact transformations under a single unifying principle and clarifies how algebraic relations between generators and transformed operators determine the polynomial degree of the transformation. We illustrate this framework for unitary coupled-cluster and involutory generators, identifying cases in which a single commutator suffices. The result establishes clear algebraic criteria for when exact unitary transformations are possible and provides new strategies for reducing their computational cost in quantum simulation.
@@ I Introduction
Unitary transformations of Hamiltonians,
@@
are fundamental tools across quantum many-body theory and quantum algorithms. They enable systematic changes of representation that can simplify correlations or entanglement in the eigenstates of a Hamiltonian. In electronic structure theory, canonical transformations are used to approximately block-diagonalize Hamiltonians before truncating the Hilbert space, facilitating downfolding approaches.[1] In variational quantum eigensolvers (VQE)[2], unitary transformations reduce circuit depth for ground- and excited-state preparation and support efficient circuit cutting in distributed implementations.[3] Quantum-inspired classical techniques, such as iterative qubit coupled cluster (iQCC)[4] and Pauli propagation methods,[5], likewise rely on Hamiltonians transformed by simple unitaries in the Heisenberg picture.
@@
in which case the transformation requires only a finite number of terms. Many important examples of such algebraic closure are known, including fermionic orbital rotations,[7] Bogoliubov transformations for both fermions and bosons,[8] and Pauli-product unitaries.[4] Comprehensive treatments of these examples can be found in Ref. 9. Recent developments in quantum computing have extended this list to linear combinations of anti-commuting Pauli operators[10] and unitary coupled-cluster (UCC) generators.[11, 12, 13]
@@ II.1 Algebraic closure from the Cayley–Hamilton theorem
where $p_{d_{V}-1}$ is a polynomial of degree $d_{V}-1$. This relation directly yields the algebraic closure condition in Eq. (5) and ensures that the Baker–Campbell–Hausdorff (BCH) series terminates.
@@ II.2 Lie–algebraic representation
$\hat{U}\hat{O}_{\mu}\hat{U}^{\dagger}=\sum_{\nu=1}^{N}[U]_{\nu\mu}\hat{O}_{\nu},$ (16)
where $U]_{\nu\mu}$ are the matrix elements of $\exp[\text{ad}_{i\theta\hat{G}}$]. To compute $U$, one can apply the BCH expansion from Eq. (3). Because the commutators close within the Lie algebra, the series can be resummed to a closed form. A faithful matrix representation $\rho:\mathfrak{L}\to\mathfrak{gl}(d_{\mathfrak{L}})$, which preserves commutation relations and has a trivial kernel, allows this resummation explicitly:
@@ II.3 Eigen–subspace projector representation
and the matrix representation of $\text{ad}_{i\hat{G}}$ has eigenvalues $\{\,i(g_{j}-g_{k})\,\}_{j,k=1}^{L}$ on the blocks $\hat{P}_{j}(\cdot)\hat{P}_{k}$. The representation space has dimension $L^{2}$, but the degree of the characteristic (and minimal) polynomial of $\text{ad}_{i\hat{G}}$ is governed only by the number of distinct nonzero differences $g_{j}-g_{k}$. Thus, in the worst case (all differences distinct and nonzero on some block) the upper bound is L(L1)L(L-1), and degeneracies or vanishing blocks reduce this degree.
@@
We define the set of relevant (nonzero) differences supported by $\hat{O}$,
@@
with the generalized Vandermonde matrix $W\in\mathbb{C}^{|S|\times(d+1)}$ defined by
[…]
@@
$\mathrm{poly}_{\mathcal{D}}^{(\Delta)}(x)\;=\;\prod_{\begin{subarray}{c}\delta\in\mathcal{D}\\ \delta\neq\Delta\end{subarray}}\frac{x-\delta}{\Delta-\delta}$ (34)
@@
If $\big\|\mathrm{poly}_{\mathcal{D}}^{(\Delta)}\!\left(\text{ad}_{\hat{G}}\right)(\hat{O})\big\|\neq 0$, then at least one block in $\mathcal{B}_{\Delta}$ is nonvanishing. Repeating this test over all $\Delta\in\mathcal{D}$ yields
$S\;=\;\big\{\,\Delta\in\mathcal{D}\;:\;\big\|\mathrm{poly}_{\mathcal{D}}^{(\Delta)}\!\left(\text{ad}_{\hat{G}}\right)(\hat{O})\big\|\neq 0\,\big\}.$ (36)
@@ II.4 Block reduction due to additional algebraic structures
Exploiting additional algebraic relations—such as commutation or anticommutation between $\hat{G}$ and $\hat{O}$—provides a direct route to identifying block sparsity without explicitly constructing the Lagrange projectors.
@@ III.1.1 Orbital rotations
$\rho(\hat{a}_{p}^{\dagger}\hat{a}_{q})]_{ab}=\delta_{ap}\,\delta_{bq}.$ (44)
In this representation, $\rho(\hat{V})=e^{iM}$. Solving Eq. (18) in this basis yields
@@ III.1.2 Heisenberg Lie algebra
$\text{ad}_{i\hat{G}}(\hat{x}_{j})=i\hbar v_{j}\hat{I},\qquad\text{ad}_{i\hat{G}}(\hat{p}_{j})=i\hbar u_{j}\hat{I}$, (53)
@@ III.2.1 Single Pauli product
For a generic $\hat{H}$, the projected blocks need not vanish; thus $S=\{0,\pm 2\}$ (recall $S$ denotes the set of distinct nonzero eigenvalue differences on nonvanishing projected blocks). From Eq. (27), the transformed Hamiltonian can be written as
@@ III.2.3 Unitary Coupled Cluster Generators
where $\hat{T}_{\alpha}=\hat{a}_{C}^{\dagger}\hat{a}_{D}$ is a single fermionic string. Let $\hat{P}_{\pm},\hat{P}_{0}$ denote the projectors onto the eigenspaces $\mathcal{V}_{\pm},\mathcal{V}_{0}$, corresponding to eigenvalues $\pm 1,0$ of $\hat{G}$, respectively. In Appendix B, we show that for any $\hat{T}_{\alpha}$, either the set $\{\hat{P}_{\pm}\hat{H}_{\alpha}\hat{P}_{\pm},\;\hat{P}_{\pm}\hat{H}_{\alpha}\hat{P}_{\mp}\}$ or the set $\{\hat{P}_{\pm}\hat{H}_{\alpha}\hat{P}_{0},\;\hat{P}_{0}\hat{H}_{\alpha}\hat{P}_{\pm}\}$ is potentially nonvanishing, but not both simultaneously.
@@
These cases agree with findings of Ref. [13]. In addition, we identify two further cases in which a single commutator suffices. In one case, the blocks $\hat{P}_{\pm}\hat{H}_{\alpha}\hat{P}_{\pm}$, $\hat{P}_{\pm}\hat{H}_{\alpha}\hat{P}_{\mp}$, and $\hat{P}_{0}\hat{H}_{\alpha}\hat{P}_{0}$ vanish, leading to $S=\{\pm 1\}$. The transformed fragment then has the form of Eq. (60), with coefficients from Eq. (31):
@@
where $\tilde{T}_{G},\tilde{T}_{\alpha}$ act on the shared spin-orbital indices $(A\cup B)\cap(C\cup D)$, while $\tilde{L}_{G}$ and $\tilde{L}_{\alpha}$ act on the disjoint sets $(A\cup B)\setminus(C\cup D)$ and $(C\cup D)\setminus(A\cup B)$, respectively. Conditions on $\{\tilde{T}_{G},\tilde{L}_{G},\tilde{T}_{\alpha},\tilde{L}_{\alpha}\}$ that produce the various nonvanishing block structures described above are derived in Appendix B. The cases with examples are summarized in Table 1, and Fig. 1 illustrates their projected-block patterns.
@@
IV Conclusions

### v2

@@ Abstract
Exact unitary transformations play a central role in the analysis and simulation of many-body quantum systems, yet the conditions under which they can be carried out exactly and efficiently remain incompletely understood. We show that exact transformations arise whenever the adjoint action of a unitary’s generator defines a linear map within a finite-dimensional operator space. In this regime, there exists a finite-degree polynomial that annihilates the adjoint map, rendering the Baker-Campbell-Hausdorff (BCH) expansion finite. We identify the role of Lie algebras and their modules in producing finite BCH expansions in all known cases. This perspective brings together previously disparate examples of exact transformations under a single unifying principle and clarifies how algebraic relations between generators and transformed operators determine the polynomial degree of the transformation. We illustrate this framework for previously known cases of efficient unitary transformations including unitary coupled-cluster and Pauli product generators. Using this framework, we propose a new class of fermionic generators that can be used for efficient transformations. The result establishes sufficient algebraic conditions for when exact unitary transformations are possible and provides new strategies for reducing their computational cost in quantum simulation and constructing feasible unitary transformations.
@@ I Introduction
Unitary transformations of a Hamiltonian,
@@
are central tools in quantum many-body theory and quantum algorithms. By systematically changing representation, they can simplify the structure of correlations and entanglement in the eigenstates of a Hamiltonian. In electronic structure theory, canonical transformations are used to approximately block-diagonalize Hamiltonians before truncating the Hilbert space, enabling downfolding approaches.[1] In variational quantum eigensolver (VQE) algorithms,[2] suitably chosen unitaries reduce circuit depth for ground- and excited-state preparation and enable efficient circuit cutting in distributed implementations.[3] In fault-tolerant settings, interaction-picture transformations are employed to reduce the cost of simulation algorithms.[4] Quantum-inspired classical techniques, such as iterative qubit coupled cluster (iQCC)[5] and Pauli propagation methods,[6] likewise rely on Hamiltonians transformed by simple unitaries in the Heisenberg picture.
@@
in which case the transformation requires only a finite number of terms. Many important examples of such algebraic closure are known, including fermionic orbital rotations,[8] Bogoliubov transformations for both fermions and bosons,[9] and Pauli-product unitaries.[5] Comprehensive treatments of these examples can be found in Ref. 10. Recent developments in quantum computing have extended this list to linear combinations of anti-commuting Pauli operators[11] and unitary coupled-cluster (UCC) generators.[12, 13, 14]
@@ II.1 Algebraic closure from the Cayley–Hamilton theorem
where $p_{d_{V}-1}$ is a polynomial of degree $d_{V}-1$. This relation directly yields the algebraic closure condition in Eq. (5) and ensures that the BCH series terminates.
@@ II.2 Lie–algebraic representation
$\hat{U}\hat{O}_{\mu}\hat{U}^{\dagger}=\sum_{\nu=1}^{d_{\mathfrak{L}}}[U]_{\nu\mu}\hat{O}_{\nu},$ (16)
where $U]_{\nu\mu}$ are the matrix elements of $\exp[\text{ad}_{i\theta\hat{G}}$]. To compute $U$, one can apply the BCH expansion from Eq. (3). Because the commutators close within the Lie algebra, the series can be resummed to a closed form. A faithful matrix representation $\rho:\mathfrak{L}\to\mathfrak{gl}(N)$, which preserves commutation relations and has a trivial kernel, allows this resummation explicitly:
@@ II.3 Eigen–subspace projector representation
and the matrix representation of $\text{ad}_{i\hat{G}}$ has eigenvalues $\{\,i(g_{j}-g_{k})\,\}_{j,k=1}^{L}$ on the blocks $\hat{P}_{j}(\cdot)\hat{P}_{k}$. The representation space has dimension $L^{2}$, but the degree of the characteristic (and minimal) polynomial of $\text{ad}_{i\hat{G}}$ is governed only by the number of distinct differences $g_{j}-g_{k}$. Thus, in the worst case (all differences distinct) the upper bound for the degree is L(L1)L(L-1), and degeneracies or vanishing blocks reduce this degree.
@@
We define the set of relevant differences supported by $\hat{O}$,
@@
with the generalized Vandermonde matrix $W\in\mathbb{C}^{|S|\times(d+1)}$ defined by [17]
$W]_{\Delta,m}\;=\;(i\Delta)^{m},\qquad\Delta\in S,\;m=0,\dots,d.$ (32)
@@
$\mathrm{Lag}_{\mathcal{D}}^{(\Delta)}(x)\;=\;\prod_{\begin{subarray}{c}\delta\in\mathcal{D}\\ \delta\neq\Delta\end{subarray}}\frac{x-\delta}{\Delta-\delta}$ (34)
@@
If $\big\|\mathrm{Lag}_{\mathcal{D}}^{(\Delta)}\!\left(\text{ad}_{\hat{G}}\right)(\hat{O})\big\|\neq 0$, then at least one block in $\mathcal{B}_{\Delta}$ is nonvanishing. Repeating this test over all $\Delta\in\mathcal{D}$ yields
$S\;=\;\big\{\,\Delta\in\mathcal{D}\;:\;\big\|\mathrm{Lag}_{\mathcal{D}}^{(\Delta)}\!\left(\text{ad}_{\hat{G}}\right)(\hat{O})\big\|\neq 0\,\big\}.$ (36)
We illustrate this test for Pauli product generators in Section III.2.1.
@@ II.4 Block reduction due to additional algebraic structures
Exploiting additional algebraic relations such as commutation or anticommutation between $\hat{G}$ and $\hat{O}$ provides a direct route to identifying block sparsity without explicitly constructing the Lagrange projectors.
@@ III.1.1 Orbital rotations
$\rho(\hat{a}_{p}^{\dagger}\hat{a}_{q})]_{ab}=\delta_{ap}\,\delta_{bq},$ (44)
where $\delta_{ap},\delta_{bq}$ are Kronecker delta functions. In this representation, $\rho(\hat{V})=e^{iM}$. Solving Eq. (18) in this basis yields
@@ III.1.2 Heisenberg Lie algebra
$\text{ad}_{i\hat{G}}(\hat{x}_{j})=\hbar v_{j}\hat{I},\qquad\text{ad}_{i\hat{G}}(\hat{p}_{j})=\hbar u_{j}\hat{I}$, (53)
@@ III.2.1 Single Pauli product
For a generic $\hat{H}$, the projected blocks need not vanish; thus $S=\{0,\pm 2\}$ (recall $S$ denotes the set of distinct eigenvalue differences on nonvanishing projected blocks). From Eq. (27), the transformed Hamiltonian can be written as
@@
The above analysis can be carried out with Lagrange polynomials of $\mathrm{ad}_{\hat{G}}$. Since the eigenvalues of $\hat{G}$ are $\pm 1$, $\mathcal{D}=\{-2,0,2\}$. Then
$\mathrm{Lag}_{\mathcal{D}}^{(-2)}(\mathrm{ad}_{\hat{G}})(\hat{\sigma}^{(\alpha)})=$ $\tfrac{1}{8}\mathrm{ad}_{\hat{G}}^{2}(\hat{\sigma}^{(\alpha)})-\tfrac{1}{4}\mathrm{ad}_{\hat{G}}(\hat{\sigma}^{(\alpha)}),$ (62)
$\mathrm{Lag}_{\mathcal{D}}^{(0)}(\mathrm{ad}_{\hat{G}})(\hat{\sigma}^{(\alpha)})=$ $-\tfrac{1}{4}\mathrm{ad}_{\hat{G}}^{2}(\hat{\sigma}^{(\alpha)})+\hat{\sigma}^{(\alpha)},$ (63)
$\mathrm{Lag}_{\mathcal{D}}^{(2)}(\mathrm{ad}_{\hat{G}})(\hat{\sigma}^{(\alpha)})=$ $\tfrac{1}{8}\mathrm{ad}_{\hat{G}}^{2}(\hat{\sigma}^{(\alpha)})+\tfrac{1}{4}\mathrm{ad}_{\hat{G}}(\hat{\sigma}^{(\alpha)}).$ (64)
When $\hat{\sigma}^{(\alpha)}$ commutes with $\hat{G}$ ($\mathrm{ad}_{\hat{G}}(\hat{\sigma}^{(\alpha)})=0$), Eqs. (62), (64) vanish while Eq. (63) does not, leading to $S=\{0\}$. When $\hat{\sigma}^{(\alpha)}$ anti-commutes with $\hat{G}$, we have $\hat{G},\hat{\sigma}^{(\alpha)}]=2\hat{G}\hat{\sigma}^{(\alpha)}$ and Eqs. (62), (64) does not vanish while Eq. (63) vanishes due to Eq. (56). This yields $S=\{-2,2\}$.
@@ III.2.3 Unitary Coupled Cluster Generators
where $\hat{T}_{\alpha}=\hat{a}_{C}^{\dagger}\hat{a}_{D}$ is a single fermionic string. Let $\hat{P}_{\pm},\hat{P}_{0}$ denote the projectors onto the eigenspaces $\mathcal{V}_{\pm},\mathcal{V}_{0}$, corresponding to eigenvalues $\pm 1,0$ of $\hat{G}$, respectively. In Appendix C, we show that for any $\hat{T}_{\alpha}$, either the set $\{\hat{P}_{\pm}\hat{H}_{\alpha}\hat{P}_{\pm},\;\hat{P}_{\pm}\hat{H}_{\alpha}\hat{P}_{\mp}\}$ or the set $\{\hat{P}_{\pm}\hat{H}_{\alpha}\hat{P}_{0},\;\hat{P}_{0}\hat{H}_{\alpha}\hat{P}_{\pm}\}$ is potentially nonvanishing, but not both simultaneously.
@@
This reduction in projected blocks thus far remain true for the transformation of a single (non Hermitian) Fermionic excitation operator $\hat{O}_{\alpha}=\hat{T}_{\alpha}$ and does not require number preservation ($|A|=|B|$). Hence our analysis agrees with findings of Ref. [14].
In addition, for the Hermitian fragments in Eq. (76), we identify two further cases in which a single commutator suffices. In one case, the blocks $\hat{P}_{\pm}\hat{H}_{\alpha}\hat{P}_{\pm}$, $\hat{P}_{\pm}\hat{H}_{\alpha}\hat{P}_{\mp}$, and $\hat{P}_{0}\hat{H}_{\alpha}\hat{P}_{0}$ vanish, leading to $S=\{\pm 1\}$. The transformed fragment then has the form of Eq. (60), with coefficients from Eq. (31):
@@
where $\tilde{T}_{G},\tilde{T}_{\alpha}$ act on the shared spin-orbital indices $(A\cup B)\cap(C\cup D)$, while $\tilde{L}_{G}$ and $\tilde{L}_{\alpha}$ act on the disjoint sets $(A\cup B)\setminus(C\cup D)$ and $(C\cup D)\setminus(A\cup B)$, respectively. Conditions on $\{\tilde{T}_{G},\tilde{L}_{G},\tilde{T}_{\alpha},\tilde{L}_{\alpha}\}$ that produce the various nonvanishing block structures described above are derived in Appendix C. The cases with examples are summarized in Table 1, and Fig. 1 illustrates their projected-block patterns.
@@
IV New many-body transformations
The proposed framework not only encompasses all the known exact transformations but also enables us to create a new class of exact transformations. This new class is motivated by the desire to combine strengths of Pauli product based generators (lowest number of eigenvalues) with the electron number preserving character of the fermionic generators. Indeed, neither a single Pauli product nor a sum of mutually anticommuting Pauli products preserves the fermionic number symmetry of the transformed Hamiltonian. We introduce fermionic, electron number preserving, generators
$\hat{G}=\hat{V}\hat{p}\hat{V}^{\dagger},$ (81)
where $\hat{V}$ is an orbital rotation [Eq. (41)] and
$\hat{p}=c_{0}+\sum_{p}c_{p}\hat{n}_{p}+\sum_{pq}c_{pq}\hat{n}_{p}\hat{n}_{q}+\cdots$ (82)
is a polynomial in the occupation operators $\hat{n}_{p}=\hat{a}_{p}^{\dagger}\hat{a}_{p}$ with only a few distinct eigenvalues. This form is motivated by the ease of constructing polynomials with a fixed number of eigenvalues from the simplest hermitian fermionic projectors, $\{\hat{n}_{p}\}$. The orbital rotations $\hat{V}$ do not alter either the eigenvalues determined by $\hat{p}$ or fermionic rank of $\hat{p}$. The form of $\hat{p}$ and $\hat{V}$ can be chosen to satisfy other desired properties, such as preservation of spin symmetries. Interestingly, a UCC generator cannot be expressed in the form of Eq. (81). [^1] A UCC generator is a linear combination of commuting Pauli products (under any fermion-qubit mapping). Hence, it can be diagonalized by a Clifford unitary rather than by an orbital rotation.
There are several approaches for constructing operators $\hat{p}$ with the target eigenvalues $g_{i}\in\Lambda$. A simple approach is to construct orthogonal projectors $\hat{P}_{i}$ as products of $\{\hat{n}_{p},1-\hat{n}_{p}\}$. The generator is then constructed as a sum of projectors with the eigenvalues as coefficients. For example, to construct a fermionic rank $2$ generator with four eigenvalues $\{g_{1-4}\}$, the projectors are constructed as $\{\hat{n}_{1}\hat{n}_{2},\hat{n}_{1}(1-\hat{n}_{2}),(1-\hat{n}_{1})\hat{n}_{2},(1-\hat{n}_{1})(1-\hat{n}_{2})\}$ and the generator takes the form
$\hat{p}$= $g_{1}\hat{n}_{1}\hat{n}_{2}+g_{2}\hat{n}_{1}(1-\hat{n}_{2})$
$+$ $g_{3}(1-\hat{n}_{1})\hat{n}_{2}+g_{4}(1-\hat{n}_{1})(1-\hat{n}_{2}).$ (83)
When the number of distinct spin-orbital indices is greater than the desired fermionic rank of the generator, some of the constructed projectors are no longer orthogonal. In such cases, this method cannot be used.
Alternatively, one can solve for coefficients $\{c_{p\cdots q}\}$ in $\hat{p}$ that satisfy the characteristic equation $p_{\Lambda}(\hat{p})=\hat{0}$. Here $p_{\Lambda}$ is the characteristic polynomial of the generator, and is defined as
$p_{\Lambda}(x)=\prod_{i}(x-g_{i}).$ (84)
The constraint $p_{\Lambda}(\hat{p})=\hat{0}$ yields a system of nonlinear equations in the coefficients, which can be solved analytically. Since non-trivial $\hat{p}$ can satisfy factors $p_{\Lambda^{\prime}}(\hat{p})=\hat{0}$ for $\Lambda^{\prime}\subset\Lambda,|\Lambda^{\prime}|\geq 2$, such solutions do not exhibit the spectrum $\Lambda$ and must be discarded.
IV.1 Fermionic reflections
To construct up to two-electron generators (fermionic rank $2$) with eigenvalues of $\pm 1$, the only operators $\hat{p}$ (up to index relabeling and a global phase of $\pm 1$) are
$\hat{p}_{1}$ $=1-2\hat{n}_{p}$
$\hat{p}_{2}$ $=1-2\hat{n}_{p}\hat{n}_{q}$
$\hat{p}_{3}$ $=1-2\hat{n}_{p}+2\hat{n}_{p}\hat{n}_{q}$
$\hat{p}_{4}$ $=1-2\hat{n}_{p}-2\hat{n}_{q}+2\hat{n}_{p}\hat{n}_{q}$
$\hat{p}_{5}$ $=1-2\hat{n}_{p}-2\hat{n}_{q}+4\hat{n}_{p}\hat{n}_{q}$
$\hat{p}_{6}$ $=1-2\hat{n}_{p}-2\hat{n}_{q}\hat{n}_{r}+2\hat{n}_{p}\hat{n}_{r}$
$\hat{p}_{7}$ $=1-2\hat{n}_{p}-2\hat{n}_{q}\hat{n}_{r}+2\hat{n}_{p}\hat{n}_{r}+2\hat{n}_{p}\hat{n}_{q}$
$\hat{p}_{8}$ $=1-2\hat{n}_{p}-2\hat{n}_{q}+2\hat{n}_{p}\hat{n}_{r}+2\hat{n}_{p}\hat{n}_{q}$
$\hat{p}_{9}$ $=1-2\hat{n}_{p}-2\hat{n}_{q}-2\hat{n}_{r}+2\hat{n}_{q}\hat{n}_{r}$
$+2\hat{n}_{p}\hat{n}_{r}+2\hat{n}_{p}\hat{n}_{q}$
$\hat{p}_{10}$ $=1-2\hat{n}_{p}-2\hat{n}_{q}-2\hat{n}_{r}\hat{n}_{s}+2\hat{n}_{p}\hat{n}_{r}$
$+2\hat{n}_{q}\hat{n}_{s}+2\hat{n}_{p}\hat{n}_{q}.$ (85)
These solutions are obtained by imposing $(c_{0}+\sum_{p}c_{p}\hat{n}_{p}+\sum_{pq}c_{pq}\hat{n}_{p}\hat{n}_{q})^{2}-\hat{1}=0$ and solving exhaustively for nontrivial values of $c_{0},\{c_{p}\},\{c_{pq}\}$. We find that $c_{0}=1$, $c_{p}\in\{0,-2\}$ and $c_{pq}\in\{0,\pm 2,4\}$, and that there are no solutions involving $5$ distinct spin-orbital indices. Thus, no solutions exist with more than four distinct spin-orbital indices. The $10$ polynomials form disjoint classes of generators because orbital rotations cannot map one of these polynomials to another (see Appendix D for a proof). Since the constructed generators are reflective ($\hat{G}^{2}=\hat{1}$), we refer to these generators as fermionic reflections.
By selecting the polynomial $\hat{p}$ and matrix $M$ defining the unitary rotations, one can optimize the generator for a specific cost function defined on the transformed Hamiltonian. One such cost function is the electronic energy of a product state
[…]
where $\hat{G}=\hat{V}\hat{p}_{j}\hat{V}^{\dagger}$. For each $j$, the matrix $M$ can be optimized classically, using the exact transformation, to determine the generator that yields the largest energy lowering. To minimize the number of terms in the transformed Hamiltonian, $M$ can instead be optimized to reduce the $\ell_{0}$-norm of the coefficients of the transformed Hamiltonian. In practice, this can be approximated by minimizing the computationally less demanding $\ell_{1}$-norm.
V Conclusions